OpenStudy (anonymous):

Find an equation of the circle that satisfies the stated conditions. (Give your answer in standard notation.) center (C)=(-2,3) passing through P(3,1)

OpenStudy (jdoe0001):

so... do you know the equation for a circle?

OpenStudy (jdoe0001):

so how to find the radius? well, the radius is the distance between the center and a point it "passes through" and you have those 2 points, how far apart are there? well \(\bf \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (anonymous):

I worked it out and im sure I did it right, but the standard notation part is getting me..

OpenStudy (jdoe0001):

the (h, k) is just the center, so your center is at (-2, 3) so \(\bf (x-h)^2+(y-k)^2=r^2 \implies (x-(-2))^2+(y-(3))^2=r^2\\ \implies (x+2)^2+(y-3)^2=r^2\)

OpenStudy (anonymous):

i had something similar but neither are right. i wonder if it's because of the standard notation